OpenStudy (anonymous):

find zeros of y=1+2cos(2x) interval; -2pi/3<=x<=2pi/3

OpenStudy (anonymous):

cos 2x =-1/2 2x=(2n+1)\(\pi \pm \frac{\pi}{3}\)

OpenStudy (anonymous):

where does the (2n+1) come from?